OpenStudy (anonymous):

Determine the nature of the solution if the equation y^2= 2/3 y+ 3/7 the y is with the whole 2/3 not just the 3

OpenStudy (anonymous):

y^2 -(2/3)y -(3/7) =0 (y - (2/6) )^2 -(3/7) = 4/36

OpenStudy (anonymous):

(y-(2/3) )^2 = (1/9) +(3/7) real solutions

OpenStudy (anonymous):

one real solutions two non real solutions two real solutions

jhonyy9 (jhonyy9):

you dont like learning ?

OpenStudy (anonymous):

I do like to learn but I feel like we were getting no where and I can understand you I am sorry

jhonyy9 (jhonyy9):

21y*2=14y +9

jhonyy9 (jhonyy9):

this is clearly now ?

OpenStudy (anonymous):

a little I just need A B C, You are kinda of backwards on your typing and something I can not understand I am very sorry. What country are you from?

OpenStudy (anonymous):

sometime not something sry*

jhonyy9 (jhonyy9):

21y*2 -14y -9=0 14 +/- radical (196 + 756) x_1,_2 = ------------------------ 42

OpenStudy (anonymous):

So two real solutions

jhonyy9 (jhonyy9):

can you continue ?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

So I would say that is two real solutions

jhonyy9 (jhonyy9):

14+/- radical 952 x_1,_2 = ----------------- can you do it ? 42

OpenStudy (anonymous):

no!!!

jhonyy9 (jhonyy9):

how do you can writing radical 952 ?

OpenStudy (anonymous):

I have tried and I have y=7+√238 7-√238 -------- , -------- 21 21

OpenStudy (anonymous):

√952

jhonyy9 (jhonyy9):

why y ?

jhonyy9 (jhonyy9):

ok sorry

jhonyy9 (jhonyy9):

so how you have wrote radical 952 ?

OpenStudy (anonymous):

what do u mean? alt+ 251 and that gives u √

jhonyy9 (jhonyy9):

and radical 238 not can being write again for we can simplify some numbers ?

OpenStudy (anonymous):

ok! that is the numbers that I got wouldnt it just be easier if I just got A, B, C please so I can go on I am on a time limit here please

jhonyy9 (jhonyy9):

yes this is right those are two real solutions what you have wrote there can we continue ?